OpenStudy (anonymous):

I give medals! Please help! There are 52 blades of grass in 2 in^2 of lawn. There are 580 blades of grass in 8 in^2 of the same lawn. Which equation models the y blades of grass found in x in^2 of lawn? y= 88x + 124 y= -88x - 124 y= -88x + 124 y= 88x - 124

OpenStudy (mertsj):

You have two points given. Find the slope of the line.

hartnn (hartnn):

could you find the slope ? take these points x1,y1= 2,52 x2,y2=8,580 then use the formula \(\large m=\dfrac{y_2-y_1}{x_2-x_1}\)

OpenStudy (anonymous):

Okay, what do I do after that?

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

then you use point slope form of equation of lines, y-y1 = m (x-x1) where x1,y1 are same, 2,52 and m you had found out to be .. ?

OpenStudy (anonymous):

m= 11.44 right?

hartnn (hartnn):

umm...nopes m= (580-52)/(8-2) =... ? you did this, right ?

OpenStudy (anonymous):

OH, sorry. 88?

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

yes, m= 88 and could you find equation of line then ?

hartnn (hartnn):

sorry for late replies.

OpenStudy (phi):

did you finish ? hartnn told you y-y1 = m (x-x1) where x1,y1 are same, 2,52 you know m, you know x1 and y1 (from (2,52) ) replace those variables with numbers what do you get ?